A bottling plant fills 1,200 bottles every two hours. The lead time is 50 minutes and...

  1. A bottling plant fills 1,200 bottles every two hours. The lead time is 50 minutes and a container accommodates 90 bottles. The safety stock is 10 percent of expected demand. How many kanban cards are needed? (Round up your answer to the next whole number.)
Answer #1

Given : d =1200 bottles every two hours =600 bottles per hour

L=50 minutes =50/60 =5/6 hour

S=10 % of expected demand=0.1*600*5/6 =50 bottles

C=90 bottles

Number of Kanban cards ,N=(d*L)+S/C

=(500+50)/90 =550/90

=6.111=7 (round up)

Number of Kanban cards needed=7